# Prompting Large Language Models for Recommender Systems: A Comprehensive Framework and Empirical Analysis

### v1

@@ Abstract.
Recently, large language models such as ChatGPT have showcased remarkable abilities in solving general tasks, demonstrating the potential for applications in recommender systems. To assess how effectively LLMs can be used in recommendation tasks, our study primarily focuses on employing LLMs as recommender systems through prompting engineering. We propose a general framework for utilizing LLMs in recommendation tasks, focusing on the capabilities of LLMs as recommenders. To conduct our analysis, we formalize the input of LLMs for recommendation into natural language prompts with two key aspects, and explain how our framework can be generalized to various recommendation scenarios. As for the use of LLMs as recommenders, we analyze the impact of public availability, tuning strategies, model architecture, parameter scale, and context length on recommendation results based on the classification of LLMs. As for prompt engineering, we further analyze the impact of four important components of prompts, i.e., task descriptions, user interest modeling, candidate items construction and prompting strategies. In each section, we first define and categorize concepts in line with the existing literature. Then, we propose inspiring research questions followed by experiments to systematically analyze the impact of different factors on two public datasets. Finally, we summarize promising directions to shed lights on future research.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXX.XXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
In order to alleviate the problem of information overload (Ni et al., 2019; Hou et al., 2022), recommender systems explore the needs of users and provide them with recommendations based on their historical interactions, which are widely studied in both industry and academia (Rendle et al., 2009; Guo et al., 2017; He et al., 2017, 2020). Over the past decade, various recommendation algorithms have been proposed to solve recommendation tasks by capturing the personalized interaction patterns from user behaviors (Kang and McAuley, 2018; Zhou et al., 2020). Despite the progress of conventional recommenders, the performance is highly dependent on the limited training data from a few datasets and domains, and there are two major drawbacks. On the one hand, traditional models lack the general world knowledge beyond interaction sequences. For complex scenarios that need to think or plan, existing methods do not have commonsense knowledge to solve such tasks (Wei et al., 2023; Harte et al., 2023; Xi et al., 2023; Liu et al., 2023a). On the other hand, traditional models cannot truly understand intentions and preferences of users. The recommendation results do not have explainability, and requirements expressed by users in explicit forms such as natural languages are difficult to consider (Li et al., 2023e; Yuan et al., 2023a; Li et al., 2023c).
@@
Recently, Large Language Models (LLMs) such as ChatGPT have demonstrated impressive abilities in solving general tasks (Haleem et al., 2022; Wu et al., 2023a), showing their potential in developing next-generation recommender systems. The advantages of incorporating LLMs into recommendation tasks are two-fold. Firstly, the excellent performance of LLMs in complex reasoning tasks indicates the rich world knowledge and superior inference ability, which can effectively compensate for the local knowledge of traditional recommenders (Sanner et al., 2023; Mysore et al., 2023; Agrawal et al., 2023). Secondly, the language modeling abilities of LLMs can seamlessly integrate massive textual data, enabling them to extract features beyond IDs and even understand user preferences explicitly (He et al., 2023; Li et al., 2023g). Therefore, researchers have attempted to leverage LLMs for recommendation tasks.
Typically, there are three ways to employ LLMs to make recommendations: (1) LLMs can serve as the recommender to make recommendation decisions, encompassing both discriminative and generative recommendations (Hou et al., 2023; Zhang et al., 2023f; Dai et al., 2023a; Gao et al., 2023; Bao et al., 2023b). (2) LLMs can be leveraged to enhance traditional recommendation models by extracting semantic representations of users and items from text corpora. The extensive semantic information and robust planning capabilities of LLMs are integrated into traditional models (Hou et al., 2022; Du et al., 2023; Wang et al., 2023a; Harte et al., 2023; Agrawal et al., 2023; Wei et al., 2023; Xi et al., 2023; Liu et al., 2023a). (3) LLMs are utilized as the recommendation simulator to execute external generative agents in the recommendation process, where users and items may be empowered by LLMs to stimulate the virtual environment (Wang et al., 2023i, g; Zhang et al., 2023e, c; Friedman et al., 2023). We mainly focus on the first scenario in this paper.
@@
Considering the gap between the general knowledge from large language models and the domain knowledge from recommendation models (Zhang et al., 2023b; Bao et al., 2023b), there are two key factors for prompting LLMs as recommenders, i.e., how to select a LLM as the foundation model and how to construct a prompt as the prompting text. As for LLMs, a growing number of open-source and closed-source models have emerged, and the same model also has different variants due to settings such as parameter scales and context lengths (Zhao et al., 2023). There are notable variations in the performance of different LLMs when it comes to general language tasks such as generation and reasoning (Wu et al., 2023a; Haleem et al., 2022). However, the performance differences of LLMs in recommendation tasks have not been fully explored. It is worth discussing how to select corresponding LLMs for specific scenarios and develop corresponding training strategies. As for the prompt, it is an important medium for interactions between humans and language models, and a well-designed prompt can better stimulate the powerful capabilities of LLMs (Liu et al., 2023f; Le Scao and Rush, 2021). To stimulate the recommendation ability of language models, prompt engineering should involve not only task description and prompting strategies for general tasks, but also the incorporation of user interest modeling and the creation of candidate items in recommender systems (Yao et al., 2023; Fan et al., 2023; Liu et al., 2023f).
@@
Table 1. An overview of the primary discoveries presented in our work. We summarize new findings in the second column as “new findings”, and conduct experiments to verify findings discussed in existing literature as “re-validated findings”.
Aspect New findings Re-validated findings
@@
LLMs $\bullet$ LLMs have cold-start recommendation capabilities, but are inferior to fully-trained recommenders. Fine-tuning LLMs can surpass traditional models. $\bullet$ The larger the parameter scale, the better the recommendation ability, while a longer maximum context length leads to worse recommendations. $\bullet$ Fine-tuning all parameters of LLMs for recommendation is more effective than parameter-efficient fine-tuning, but more training time is required. $\bullet$ Instruction tuning can enhance the fine-tuning results of LLMs on recommendations (Bao et al., 2023b; Touvron et al., 2023b). $\bullet$ In few-shot training scenarios, LLMs are more capable of adapting to recommendation tasks compared to traditional models (Liao et al., 2023; Fu et al., 2023). $\bullet$ Despite capabilities, there are limitations leveraging LLMs as recommenders, such as position bias (Hou et al., 2023; Ma et al., 2023) and lack of domain knowledge (Yao et al., 2023; Zheng et al., 2023).
Prompts Task description
$\bullet$ Our framework can be adapted to point-wise, pair-wise and list-wise recommendation tasks. $\bullet$ Different recommendation tasks can be implemented by LLMs through prompts (Wang et al., 2023c; Geng et al., 2022).
User interest Modeling
@@
$\bullet$ For short-term interest, it is preferable to summarize recent interactions into text and recommend. $\bullet$ For long-term interest, it is useful to maintain a personalized memory to store and retrieve. $\bullet$ Long-term preferences and short-term intentions should be effectively combined to model users. $\bullet$ For short-term interest, only truncating the most recent items is not the optimal strategy (Lin et al., 2023b). $\bullet$ Increasing the number of historical items to represent users brings insignificant gains for LLMs (Hou et al., 2023). $\bullet$ Personalized user profiles and customized item descriptions assist in the user interest modeling for LLM-based recommendations (Yao et al., 2023; Shu et al., 2023).
@@
$\bullet$ Candidate items construction for LLMs is crucial to the final recommendation results. $\bullet$ Retrieving candidate items by traditional recommendation models first, and then re-ranking items by LLMs can further improve the results, but the performance varies on specific methods and datasets. $\bullet$ LLMs can select from a given candidate item list (Dai et al., 2023a; Liu et al., 2023c), as well as directly generate recommendation results (Liu et al., 2023a; Wang et al., 2023e). $\bullet$ Indexing methods and grounding strategies of items are important factors affecting the effectiveness of LLM-based recommendations (Lin et al., 2023c; Liao et al., 2023; Li et al., 2023b).
@@
$\bullet$ In chain-of-thought prompting, specific problem decomposition is required for recommendation tasks rather than general prompts. $\bullet$ The few-shot prompting strategy has insignificant advantage in recommendation scenarios. $\bullet$ Although provided with historical items in chronological order, LLMs still needs explicit guidance to understand the importance of recent items (Hou et al., 2023; Ma et al., 2023). $\bullet$ Role-playing and expert-like prompts can leverage the capabilities of LLMs in specific fields (Jin et al., 2023; Wang et al., 2023h).
Overall $\bullet$ Leveraging LLMs as recommenders lies in stimulating the general knowledge of LLMs and integrating the domain knowledge with the user interest. $\bullet$ LLMs are restricted by the unacceptable inference time (Li et al., 2023b), expensive memory cost (Touvron et al., 2023b), limited context length (Lin et al., 2023b) and black-box abilities (Li et al., 2023h).
@@
In order to conduct our analysis for ProLLM4Rec, we formalize the input of LLMs for recommendation into natural language prompts with two key aspects: LLMs and prompts, and explain how our framework can be generalized to various recommendation scenarios and tasks. As for the use of LLMs as recommenders, we analyze the impact of the public availability, tuning strategies, model architecture, parameter scale, and context length on recommendation results based on the classification of LLMs. As for prompt engineering, we further analyze the impact of four important components of prompts, i.e., task description, user interest modeling, candidate items construction, and prompting strategies. Given personalized prompts that include task description and user interest, the LLM selects, generates, or explains candidate items based on general world knowledge and personalized user profiles. For each module, we first define and categorize concepts in line with the existing literature. Then, we propose inspiring research questions, followed by detailed experiments to systematically analyze the impact of different conditions on the recommendation performance. Based on the empirical analysis, we finally summarize empirical findings for future research.
@@
  • We derive a general framework ProLLM4Rec to sum up existing work of utilizing LLMs as foundation models for recommendation, which can be generalized to multiple scenarios and tasks by different LLMs and prompts.
  • We provide a systematic analysis on leveraging LLMs for recommendation, focusing on two aspects: LLMs and prompt engineering. The use of LLMs includes analysis of public availability, tuning strategies, model architecture, parameter scale, and context length. Moreover, prompt engineering consists of discussions on task description, user interest modeling, candidate items construction and prompting strategies. For each aspect, we define and describe with concepts first, and then provide reference solutions with experiments.
@@
In what follows, we first review the related work in Section 2. In Section 3, we present our proposed general framework and its instantiation, and introduce overall settings of the following experiments. As the core components of this paper, we discuss two main aspects of ProLLM4Rec, i.e., LLMs and prompts in Section 4 and Section 5, respectively. For each aspect, we generalize key factors that affect recommendation results, and conduct corresponding experiments to summarize empirical findings. At last, Section 6 concludes this paper and sheds lights on future directions.
@@ 2.1. Recommender Systems
For tackling the challenge of information overload (Ni et al., 2019; Thorat et al., 2015; Yuan et al., 2023b), recommender systems have become pivotal tools for delivering personalized contents for users across various domains. In line with previous studies, recommendation algorithms aim to derive user preferences and behavioral patterns from their historical interactions. The most common technique for the interaction-based recommendation is Collaborative Filtering (CF) (Sarwar et al., 2001; Su and Khoshgoftaar, 2009), which recommends items based on preferences of similar users. Matrix Factorization (MF) (Koren et al., 2009) is a prevalent approach in collaborative filtering, and it constructs embedding representations for users and items from the interaction matrix, facilitating the algorithm to calculate similarity scores efficiently. Furthermore, Neural Collaborative Filtering (NCF) (He et al., 2017), integrating deep neural networks, replaces the inner product used in MF with a neural architecture, thereby demonstrating better performance than previous methods. Contemporary advancements in deep neural network architectures have enhanced the integration of user and item embeddings (Liu et al., 2023b). For example, since recommendation data can be represented as graph-structured data, Graph Neural Network (GNN) (Wu et al., 2022) can be utilized to encode the information of the interaction graph (nodes consist of users and items), and generate meaningful representations via message propagation and contrastive learning strategies (Wang et al., 2019; He et al., 2020; Wu et al., 2021a; Lin et al., 2022). As Pre-trained Language Models (PLM) gain prominence, there is a growing interest in pre-trained large-scale recommendation models powered by PLMs (Zhou et al., 2020; Zhang et al., 2021; Hou et al., 2022). In addition to user-item pairs and IDs, content-based recommendation algorithms leverage auxiliary modalities such as textual and visual information to augment user and item representations in recommendation tasks (Qiu et al., 2021; Yuan et al., 2023b; Wang, 2023).
@@ 2.2. Large Language Models for Recommender Systems
Large Language Models (LLMs) are a cutting-edge advancement in artificial intelligence that excel in understanding and generating human-like texts (Touvron et al., 2023a; Zeng et al., 2022; Raffel et al., 2020). LLMs are usually transformer-based models and trained on vast amounts of textual data with billions of parameters, allowing them to comprehend contexts, generate coherent sentences, and even mimic human conversations (Haleem et al., 2022; Wu et al., 2023a). Through this process, LLMs have shown prominent potentials in the field of Natural Language Processing (NLP), and have demonstrated various incredible capabilities in dealing with complex NLP tasks, including but not limited to In-Context Learning (ICL) (Brown et al., 2020), instruction following (Touvron et al., 2023b) and step-by-step reasoning abilities (Zhao et al., 2023). Recently, LLMs have been increasingly integrated into recommender systems to provide personalized recommendations (Li et al., 2023l; Wu et al., 2023b). Recent studies have explored the fusion of LLMs with recommender systems, which can be divided into the three paradigms, i.e., LLM as recommendation model (Section 2.2.1), LLM improves recommendation models (Section 2.2.2) and LLM as recommendation simulator (Section 2.2.3) as follows.
2.2.1. LLM as Recommendation Model

### v2

# Tapping the Potential of Large Language Models as Recommender Systems: A Comprehensive Framework and Empirical Analysis

@@ Abstract.
Recently, Large Language Models (LLMs) such as ChatGPT have showcased remarkable abilities in solving general tasks, demonstrating the potential for applications in recommender systems. To assess how effectively LLMs can be used in recommendation tasks, our study primarily focuses on employing LLMs as recommender systems through prompting engineering. We propose a general framework for utilizing LLMs in recommendation tasks, focusing on the capabilities of LLMs as recommenders. To conduct our analysis, we formalize the input of LLMs for recommendation into natural language prompts with two key aspects, and explain how our framework can be generalized to various recommendation scenarios. As for the use of LLMs as recommenders, we analyze the impact of public availability, tuning strategies, model architecture, parameter scale, and context length on recommendation results based on the classification of LLMs. As for prompt engineering, we further analyze the impact of four important components of prompts, i.e., task descriptions, user interest modeling, candidate items construction and prompting strategies. In each section, we first define and categorize concepts in line with the existing literature. Then, we propose inspiring research questions followed by detailed experiments on two public datasets, in order to systematically analyze the impact of different factors on performance. Based on our empirical analysis, we finally summarize promising directions to shed lights on future research.
@@
copyright: acmlicensedjournalyear: 2025doi: XXX.XXXccs: Information systems Recommender systemsccs: Information systems Language models
@@ 1. Introduction
In order to alleviate the problem of information overload (Ni et al., 2019; Hou et al., 2022), recommender systems explore user needs and provide them with recommendations based on their historical interactions, which are widely studied both in industry and academia (Rendle et al., 2009; Guo et al., 2017; He et al., 2017, 2020). During the past decade, various recommendation algorithms have been proposed to solve recommendation tasks by capturing personalized interaction patterns from user behaviors (Kang and McAuley, 2018; Zhou et al., 2020). Despite the progress of conventional recommenders, the performance is highly dependent on the limited training data from a few datasets and domains, and there are two major drawbacks. On the one hand, traditional models lack the general knowledge of the world beyond interaction sequences. For complex scenarios that need to think or plan, existing methods do not have commonsense knowledge to solve such tasks (Wei et al., 2024b; Harte et al., 2023; Xi et al., 2023; Liu et al., 2023a). On the other hand, traditional models cannot truly understand intentions and preferences of users. The results of the recommendation do not have explainability, and the requirements expressed by users in explicit forms such as natural languages are difficult to consider (Li et al., 2023e; Yuan et al., 2023a; Li et al., 2023d).
@@
Recently, Large Language Models (LLMs) such as ChatGPT have demonstrated impressive abilities in solving general tasks (Haleem et al., 2022; Wu et al., 2023), showing their potential in the development of next-generation recommender systems. The advantages of incorporating LLMs into recommendation tasks are two-fold. Firstly, the excellent performance of LLMs in complex reasoning tasks indicates the rich world knowledge and superior inference ability, which can effectively compensate for the local knowledge of traditional recommenders (Sanner et al., 2023a; Mysore et al., 2023; Agrawal et al., 2023). Secondly, the language modeling abilities of LLMs can seamlessly integrate massive textual data, enabling them to extract features beyond IDs and even explicitly understand user preferences (He et al., 2023; Li et al., 2023g). Therefore, researchers have attempted to leverage LLMs for recommendation tasks. Typically, there are three ways to employ LLMs to make recommendations: (1) LLMs can serve as the recommender to make recommendation decisions, encompassing both discriminative and generative recommendations (Hou et al., 2024; Zhang et al., 2023e; Dai et al., 2023a; Gao et al., 2023; Bao et al., 2023b). (2) LLMs can be used to improve traditional recommendation models by extracting semantic representations of users and items from text corpora. The extensive semantic information and robust planning capabilities of LLMs are integrated into traditional models (Hou et al., 2022; Du et al., 2023; Wang et al., 2023a; Harte et al., 2023; Agrawal et al., 2023; Wei et al., 2024b; Xi et al., 2023; Liu et al., 2023a). (3) LLMs are used as the recommendation simulator to execute external generative agents in the recommendation process, where LLMs can empower users and items to stimulate the virtual environment (Wang et al., 2023h, e; Zhang et al., 2023d, 2024a; Friedman et al., 2023). We mainly focus on the first scenario in this paper.
@@
Considering the gap between general knowledge from LLMs and domain knowledge from recommendation models (Zhang et al., 2023b; Bao et al., 2023b), there are two key factors to leverage LLMs as recommenders, i.e., how to select an LLM as the foundation model and how to construct task-specific prompts. As for LLMs, a growing number of open-source and closed-source models have emerged, and the same model also has different variants due to settings such as parameter scales and context lengths (Zhao et al., 2023). It is worth discussing how to select corresponding LLMs for specific scenarios and develop corresponding training strategies. As for prompts, it is an important medium for interactions between humans and language models, and a well-designed prompt can better stimulate the powerful capabilities of LLMs (Liu et al., 2023e; Scao and Rush, 2021). To stimulate the recommendation ability of LLMs, prompt engineering should involve not only task description and prompting strategies for general tasks, but also the incorporation of user interest and candidate items in recommender systems (Yao et al., 2023; Zhao et al., 2024; Liu et al., 2023e).
@@
Table 1. An overview of the primary discoveries presented in our work. We summarize our findings in the second column as “our findings”, and verify findings in existing literature as “re-validated findings”.
Aspect Our Findings Re-validated Findings
@@
LLMs $\bullet$ The long-chain reasoning capabilities and slow-thinking process of LLMs are optimized for Mathematics and coding tasks, which may not be suitable for the recommendation task. $\bullet$ The larger the parameter scale, the better the ability, while a longer maximum context length leads to worse recommendations. $\bullet$ Fine-tuning all parameters of LLMs is more effective than parameter-efficient fine-tuning such as LoRA, but more training time is required. Similarly, a larger rank in LoRA means that more parameters can be fine-tuned, generally achieving better recommendations. $\bullet$ LLMs have zero-shot recommendation capabilities, but are inferior to fully-trained recommenders. Fine-tuning LLMs can surpass traditional recommendation models (Dai et al., 2023a; Hou et al., 2024). $\bullet$ Instruction tuning can enhance fine-tuning results of LLMs on recommendations (Bao et al., 2023b; Touvron et al., 2023b). $\bullet$ In few-shot training scenarios, LLMs are more capable of adapting to recommendation tasks compared to traditional models (Liao et al., 2023; Fu et al., 2023). $\bullet$ There are limitations leveraging LLMs as recommenders, such as position bias (Hou et al., 2024; Ma et al., 2023) and lack of domain knowledge (Yao et al., 2023; Zheng et al., 2024).
Prompts User Interest Modeling
@@
$\bullet$ The recommendation performance of retrieval-based and generation-based methods for user interest modeling is generally better than that of memory-based methods, indicating the importance of the interest memory. $\bullet$ Long-term interest and short-term interest can complement each other, and it is preferable to combine the two with personalized descriptions for personalized query and memory. $\bullet$ For short-term interest, only truncating the most recent items is not optimal (Lin et al., 2024a). $\bullet$ Increasing the number of historical items to represent users brings insignificant gains (Hou et al., 2024). $\bullet$ Personalized user profiles and customized item descriptions assist in user modeling for LLM-based recommendations (Yao et al., 2023; Shu et al., 2023). $\bullet$ Retrieval-enhanced items can improve sequential comprehension (Lin et al., 2024a).
@@
$\bullet$ LLMs exhibit limitations in retaining critical information regarding the candidate list from prompts, potentially leading to hallucination and responses that deviate from the task. $\bullet$ Description-based identifiers outperform token-based when grounding items. $\bullet$ Retrieving candidate items by traditional models and then re-ranking them by LLMs can further improve the results (Hou et al., 2024; Yang et al., 2023a). $\bullet$ Indexing methods and grounding strategies of items are important factors affecting the effectiveness of recommendations (Lin et al., 2023b; Liao et al., 2023; Li et al., 2023c).
@@
$\bullet$ Specific problem decomposition is required for recommendation in CoT prompting. $\bullet$ The few-shot prompting strategy has insignificant advantage in recommendations. $\bullet$ Explicit feedback is better than implicit feedback to fine-tune LLMs for recommendation. $\bullet$ Although provided with historical items in chronological order, LLMs still need explicit guidance to understand recent items (Hou et al., 2024; Ma et al., 2023). $\bullet$ Role-playing and expert-like prompts can leverage the capabilities of LLMs (Jin et al., 2023; Wang et al., 2023f).
@@
In order to conduct our analysis for LLM-RS, we formalize the input of LLMs for recommendation into natural language prompts with two key aspects: LLMs and prompts, and explain how our framework can be generalized to various recommendation scenarios and tasks. As for the use of LLMs as recommenders, we analyze the impact of the public availability, tuning strategies, model architecture, parameter scale, and context length on recommendation results based on the classification of LLMs. As for prompt engineering, we further analyze the impact of four important components of prompts, i.e., task description, user interest modeling, candidate items construction, and prompting strategies. Given personalized prompts that include task description and user interest, the LLM selects, generates, or explains candidate items based on general world knowledge and personalized user profiles. For each module, we first define and categorize concepts in line with the existing literature. Then, we propose inspiring research questions, followed by detailed experiments to analyze the impact of different conditions on the recommendation performance. Based on the empirical analysis, we finally summarize empirical findings for future research.
@@
  • We derive a general framework LLM-RS to sum up existing work of utilizing LLMs as foundation models for recommendation, which can be generalized to multiple scenarios and tasks by different LLMs and prompts.
  • We provide a systematic analysis on leveraging LLMs as recommender systems, focusing on two aspects: LLMs and prompt engineering. The use of LLMs includes analysis of public availability, tuning strategies, model architecture, parameter scale, and context length. Moreover, prompt engineering consists of discussions on task description, user interest modeling, candidate items construction and prompting strategies. For each aspect, we define and describe with concepts first, and then provide reference solutions with experiments.
@@
In what follows, we first review the related work in Section 2. In Section 3, we present our proposed general framework and its instantiation, and introduce overall settings of the following experiments. As the core components of this paper, we discuss two main aspects of LLM-RS, i.e., LLMs and prompts in Section 4 and Section 5, respectively. For each aspect, we generalize key factors that affect recommendation results, and conduct corresponding experiments to summarize empirical findings. At last, Section 6 concludes this paper and sheds lights on future directions.
@@ 2.1. Recommender Systems
For tackling the challenge of information overload (Ni et al., 2019; Yuan et al., 2023b), recommender systems have become pivotal tools for delivering personalized contents for users across various domains. Based on previous studies, research has delineated two primary categories for common recommendation methods: interaction-based recommendation and content-based recommendation (Thorat et al., 2015). In line with previous studies, recommendation algorithms aim to derive user preferences and behavioral patterns from their historical interactions. The most common technique for the interaction-based recommendation is Collaborative Filtering (CF) (Sarwar et al., 2001; Su and Khoshgoftaar, 2009), which recommends items based on preferences of similar users. Matrix Factorization (MF) (Koren et al., 2009) is a prevalent approach in collaborative filtering, and it constructs embedding representations for users and items from the interaction matrix, facilitating the algorithm to calculate similarity scores efficiently. Furthermore, Neural Collaborative Filtering (NCF) (He et al., 2017), integrating deep neural networks, replaces the inner product used in MF with a neural architecture, thereby demonstrating better performance than previous methods. Contemporary advancements in deep neural network architectures have enhanced the integration of user and item embeddings (Liu et al., 2023b). For example, since recommendation data can be represented as graph-structured data, Graph Neural Network (GNN) (Wu et al., 2022) can be utilized to encode the information of the interaction graph (nodes consist of users and items), and generate meaningful representations via message propagation and contrastive learning strategies (Wang et al., 2019; He et al., 2020; Wu et al., 2021a; Lin et al., 2022). As Pre-trained Language Models (PLM) gain prominence, there is a growing interest in pre-trained large-scale recommendation models powered by PLMs (Zhou et al., 2020; Zhang et al., 2021; Hou et al., 2022). Besides user-item pairs and IDs, content-based recommendation algorithms leverage auxiliary modalities such as textual and visual information to augment user and item representations in recommendation tasks (Qiu et al., 2021; Yuan et al., 2023b; Wang, 2024).
@@ 2.2. Large Language Models for Recommender Systems
Large Language Models (LLMs) are a cutting-edge advancement in artificial intelligence that excel in understanding and generating human-like texts (Touvron et al., 2023a; Zeng et al., 2023; Raffel et al., 2020). LLMs are usually transformer-based models and trained on vast amounts of textual data with billions of parameters, allowing them to comprehend contexts, generate coherent sentences, and even mimic human conversations (Haleem et al., 2022; Wu et al., 2023). Through this process, LLMs have shown prominent potentials in the field of Natural Language Processing (NLP), and have demonstrated various incredible capabilities in dealing with complex NLP tasks, including but not limited to In-Context Learning (ICL) (Brown et al., 2020), instruction following (Touvron et al., 2023b) and step-by-step reasoning abilities (Zhao et al., 2023). Recently, LLMs have been increasingly integrated into recommender systems to provide personalized recommendation (Li et al., 2024d; Wu et al., 2024b; Zhao et al., 2024). Recent studies have explored the fusion of LLMs with recommender systems, which can be divided into the three paradigms, i.e., LLMs as recommender systems (Section 2.2.1), LLMs improve recommender systems (Section 2.2.2) and LLMs as recommendation simulator (Section 2.2.3) as follows.
2.2.1. LLMs as Recommender Systems